OpenStudy (bold2323):

Select the equation that contains the points (-1, -8) and (3, 4). a. y-4=-2(x-3) b. y+8=-2(x+1) c. y+4=3(x+3) d. y-4=3(x-3)

TheSmartOne (thesmartone):

First find the slope. \(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\) Where the two points are in the form: \(\sf\Large (\color{purple}{x_1},\color{red}{y_1})~and~(\color{green}{x_2},\color{blue}{y_2})\)

TheSmartOne (thesmartone):

can you find the slope between (-1, -8) and (3, 4)?

OpenStudy (skullpatrol):

Once you find the slope m, recall the point-slope formula: $$\Huge y-y_0=m(x-x_0)$$